# XUE: The CO2-rich terrestrial planet-forming region of an externally irradiated Herbig disk

### v1

@@ 6 Conclusions
We present JWST/MIRI MRS observations of the inner part of the Herbig disk XUE 10 (⋆ 3 ⊙), located in the massive NGC 6357 cluster (1.69 kpc), and of estimated external FUV irradiation exposure of about 5 103 than the Solar neighborhood and nearby star-forming regions. We summarize our conclusions as follows.
@@
We report the first simultaneous detection ( 5) of four isotopic species of CO2 (12CO2, 13CO2, 16O12C18O, 16O12C17O) in a protoplanetary disk, faint CO emission (2), atomic hydrogen HI Pf (8), and an upper limit on the H2O content (tot 1018 cm-2).
@@
16O12C18O and 16O12C17O may be isotopically enhanced compared to the oxygen isotopic ratios (16O/18O and 16O/17O) found in the interstellar medium and in the Solar System, albeit line optical depth effects cannot be ruled out.
@@
We detect PAHs emission at 6.2, 7.7–8.6 and 11 m, and of silicate aggregates at 10–11 m, consistent with other observations of IMTT and Herbig disks, except for a 15.5–17 m feature whose origin remains unclear.
@@
We propose that the observed chemistry in XUE 10 depends on transport processes. Water is removed either via advection or strong photo-dissociation by stellar UV irradiation. CO2 is locally enhanced in the gas-phase via the CO + OH CO2 + H reaction, facilitated by water removal in relatively cold gas (gas 300 K). External FUV irradiation supports the observed CO2–H2O dichotomy, and also the potential detection of CO2 oxygen isotopic anomalies in the inner disk.
A controlled infrared sample of IMTT and Herbig disks both in isolated and irradiated environments is needed at JWST resolution to perform a statistical comparison of the disk properties around young intermediate-mass stars, and corroborate trends as a function of the stellar mass. Major modeling efforts are also needed to understand whether the ?line-poor? evidence in disks around intermediate-mass stars is merely an observational bias due to a limitation in sensitivity against the bright dust continuum, or whether it is symptomatic of underlying crucial physical differences between IMTT/Herbig disks and the lower-mass counterparts. The detection of several CO2 isotopologues in XUE 10 highlights the possibility of investigating the isotopic composition of the molecular carriers of the CHNOPS life elements at the protoplanetary disk stage.
@@ Appendix A Local continuum normalization
In order to fit the molecular emission in the observed spectrum, it is first needed to subtract the local dust continuum. We therefore fitted and subtracted a spline function in each wavelength window of interest selecting line-free regions, as performed e.g., by Grant et al. (2023). We evaluated the spectral uncertainty as the standard deviation of the flux in a line-free region close to each fitted range. The continuum points used to determine the spline in all windows and the estimated spectral uncertainty are reported in Figure 10. In the 12.9-17.6 m wavelength region with detected emission from the CO2 isotopologues, these species are so abundant to create a pseudo-continuum on top of the dust continuum, corresponding to their strongest emitting features, i.e., fundamental -branch and hot bandheads, particularly those of 12CO2 and 13CO2. This makes it exceptionally challenging to subtract the local dust continuum. We therefore chose the continuum points based on templates of each isotopologue emission for different values of temperature, column density and equivalent emitting radius overlaid on the observed spectrum. We then continuum-subtracted the observed spectrum before fitting each species, selecting line-free regions or regions where the contamination from the other isotopologues was minimal. Note that in Figure 10 it is only shown the interpolated continuum in the original spectrum between 12.9–17.6 m before 12CO2 fit and subtraction.
@@
The only gas-phase molecular species capable of creating such a strong feature is diacetylene (C4H2), as shown in Figure 11. In order to reproduce this feature, the gas slab modeling approach requires C4H2 to emit with a very high column density of 3.3 1019 cm-2 at 221 K, and an equivalent emitting radius of 3.3 au ( = 8.1; Figure 13). However, the optically thin line emission expected to arise on top of the optically thick pseudo-continuum is not reproduced in the spectrum, nor the C4H2 line emission which should be visible at 14–14.5 m. The overlap of an optically thick and an optically thin slab model component does not solve the issue.
One may expect to see not only this species, but a ?forest? of hydrocarbonic molecules, starting from the more common acetylene (C2H2) species. This has been observed in the carbon-rich inner regions of some disks around very low-mass stars (e.g., Tabone et al. 2023; Arabhavi et al. 2024; Long et al. 2025; Kanwar et al. 2024). The emission of C4H2 may also be produced by ongoing thermal or UV photo-destruction of carbonaceous grains in the surface layer of XUE 10 (e.g. Visser et al. 2007; Kress et al. 2010; Anderson et al. 2017).
Another possibility may be the emission of PAHs or dust aggregates never seen before in IMTT or Herbig disks at Spitzer resolution (e.g., Meeus et al. 2001; Keller et al. 2008; Valegård et al. 2021). We tested the hypothesis of dust aggregates by fitting the JWST mid-IR dust continuum with the DuCK component of the Dust Continuum Kit with Line emission from Gas (DuCKLinG) modeling tool (Kaeufer et al. 2024). We input the available dust grain opacities for Am Mg-olivine, Am Mg-pyroxene, silica, enstatite and forsterite of 0.1, 1.0, 2.0, 3.0, 4.0, and 5.0 m in size. To model the stellar component, we gave in input the stellar parameters derived for XUE 10 (see Section 3.2). For the disk, we assumed standard parameter priors for the rim, the midplane and the surface, compatible with modeling ranges and observations of Herbig disks. Nonetheless, the ?bump? is not reproduced by any combination of these common solid-state dust species. A more detailed analysis of the dust component and its mineralogy is beyond the scope of this paper, and it will be the subject of a forthcoming work.
@@ Appendix C Spectral classification
The adopted grid spans over 5000 eff 12 000 K, in effective temperature, 2.0 log(g) 6.0 in surface gravity, and -2.0 [Fe/H] +1.0 in metallicity. In order to control for dependencies from the luminosity class over the grid, we first normalized the spectra. To do so, we created a simple Python routine in which a spectrum is split in two parts, over 7730–8400 and 8400–9480 respectively, before and after the Paschen jump. Each part is individually normalized by subtracting a high-order polynomial, and then the two chunks are stitched together into the final normalized spectrum. Note that this routine has been written and optimized for the sole purpose of this analysis.
We proceeded by comparing the normalized (telluric-corrected) FORS2 spectrum with normalized PHOENIX templates, using the reduced -test of the form:
@@
In order to evaluate the uncertainty on the fitted quantities, we performed an MCMC simulation using the Python emcee package (Foreman-Mackey et al. 2013). We set up uniform priors in the adopted PHOENIX grid for the three fitted parameters (eff, log(g), [Fe/H]). We chose the initial positions of the MCMC walkers as the estimated chi-square best-fit values. We imposed the size of the random ?kick? of the walkers from their initial position according to the dynamic range of each parameter space i.e., 500 K in temperature, and 0.5 for surface gravity and metallicity, to ensure a good parameter space sampling.
@@
We run the simulation with 300 walkers and 100 000 steps, adjusted to minimize the autocorrelation times, and increase the number of effective samples in a reasonable amount of computational time. From the posterior distribution of samples, we discarded the first 20 000 samples (20% of the number of steps), and kept a sample from each walker every 20th step, i.e., thinning the posterior distribution by 20. The corner plot of the simulation is shown in Figure 12. The 1 and 2 2D contours are well-defined around the initial positions (blue square). The statistical uncertainties () on the fitted quantities are reported on top of the diagonal histograms. The median parameter values of the posterior distribution are 7000 K (eff), 4.0 (log(g)), and +0.5 ([Fe/H]), in agreement with the chi-square fit within error bars.
@@
According to our analysis, XUE 10 can be classified as a F0-F1 type PMS (Gray & Corbally 2009). We combined the spectral information derived with PHOENIX with multi-band photometry to assess visual extinction (V), age, stellar mass (⋆), and bolometric luminosity (bol) of XUE 10. The photometric data include PAN-STARRS, SDSS, Gaia, APASS, and HST in the optical range; DENIS, 2MASS, VISTA, and UKIRT in the mid-IR, and Spitzer and WISE in the far-IR range. Assuming a total-to-selective extinction factor V = 3.3 (Massi et al. 2015; Russeil et al. 2017; Fouesneau et al. 2022), we estimate with VOSA a visual extinction of V 4.8. This result is consistent with the V range of 4.6–4.8 derived using the XUE 10 Gaia color (bp-rp) where a range of spectral types A7–F1 is assumed. We then compared the PHOENIX best-fit template with PMS isochrones of varying age (0.7, 1, 3, 5, 10 Myr) at solar metallicity and at the constrained V, using the PARSEC database (Bressan et al. 2012)666http://stev.oapd.inaf.it/cgi-bin/cmd. Note that PARSEC isochrones are computed by default for an V = 3.1. However, the inferred uncertainties on eff and log(g) are too large to reliably locate XUE 10 on a specific isochrone in this range. XUE 10 is most likely consistent with a 1–3 Myr old PMS in NGC 6357. Starting from the VVV VISTA apparent -band magnitude J = 12.004 0.001, we derive a bolometric luminosity bol = 69 5 ⊙, adopting the bolometric correction in -band for a F1 PMS of Pecaut & Mamajek (2013), the constrained V ( 0.2 from Gaia), and V = 3.3. Considering a PARSEC isochrone of 3 Myr extincted by a factor 4.8 in visual magnitude, XUE 10’s stellar counterpart is constrained to a mass of 2.5–3.0 ⊙.
@@ Appendix D Fitting tests
Assuming LTE conditions, the fitting procedure consists in comparing a grid of synthetic spectra computed for different values of (gas, tot, em) with the observed continuum-subtracted spectrum through the reduced chi-square test of the form of Eq. 1 (Tabone et al. 2023). Here, obs is the number of MIRI resolution elements covering species features in each fitted wavelength window, obs and mod are respectively the observed and the model fluxes to compare, and is the flux uncertainty estimated in a line-free region closed to the spectral region where the species is modeled (see Figure 10). This approach has been extensively used for the spectral fitting of many disk sources so far (see e.g., Grant et al. 2023; Perotti et al. 2023; Gasman et al. 2023; Ramírez-Tannus et al. 2023; Schwarz et al. 2024; Temmink et al. 2024a), yielding best-fit parameters which are also consistent with the outcome of more sophisticated approaches, e.g., ProDiMo thermo-chemical models (Woitke et al. 2024b) and MCMC simulations (Temmink et al. 2024a). For all molecules, we run a grid of models varying tot (log10-space) in steps of 0.2, gas in steps of 20 K, and em (log10-space) in steps of 0.09. The dynamic ranges explored for each parameter are in Table 4. Given the limited spectral resolution of the MIRI MRS instrument (Jones et al. 2023), we convolved the model spectra with a resolution = 3550 at 4.9–5.2 m, = 3355 at 6.4–7.6 m, = 2600 between 12.9 and 17.6 m, and = 1730 at 18–25 m, where we fit the different species. Then, we resampled the convolved models on the wavelength grid of the observations.
@@ Appendix E Reduced chi-square maps
The maps for all fitted species are shown in Figure 13 in order of increasing wavelength in the spectrum. The results for 13CO2, 16O12C18O, and 16O12C17O are reported assuming the best-fit emitting radius found for 12CO2 (em = 1.15 au), and leaving em as a free parameter (see Section 4.2).

### v2

@@ 6 Conclusions
We have presented JWST/MIRI MRS observations of the inner part of the Herbig disk XUE 10 (⋆ 3 ⊙) located in the massive NGC 6357 cluster (1.69 kpc) and which has an estimated external FUV irradiation exposure of about 5 103 stronger than the solar neighborhood and nearby star-forming regions. We summarize our conclusions as follows:
@@
We report the first simultaneous detection ( 5) of four isotopic species of CO2 (12CO2, 13CO2, 16O12C18O, 16O12C17O) in a protoplanetary disk, a faint CO emission (2), atomic hydrogen HI Pf (8), and an upper limit on the H2O content (tot 1018 cm-2).
@@
The species 16O12C18O and 16O12C17O may be isotopically enhanced compared to the oxygen isotopic ratios (16O/18O and 16O/17O) found in the ISM and in the Solar System, though line optical depth effects cannot be ruled out.
@@
We detected PAH emission at 6.2, 7.7–8.6, and 11 m and of silicate aggregates at 10–11 m, which is consistent with other observations of IMTT and Herbig disks, except for a 15.5–17 m feature whose origin remains unclear.
@@
We propose that the observed chemistry in XUE 10 depends on transport processes. Water is removed either via advection or strong photo-dissociation by stellar UV irradiation. CO2 is locally enhanced in the gas phase via the CO + OH CO2 + H reaction, facilitated by water removal in relatively cold gas (gas 300 K). External FUV irradiation supports the observed CO2–H2O dichotomy and the potential detection of CO2 oxygen isotopic anomalies in the inner disk.
A controlled infrared sample of IMTT and Herbig disks both in isolated and irradiated environments is needed at JWST resolution to perform a statistical comparison of the disk properties around young intermediate-mass stars and to corroborate trends as a function of the stellar mass. Major modeling efforts are also needed to understand whether the ?line-poor? evidence in disks around intermediate-mass stars is merely an observational bias due to a limitation in sensitivity against the bright dust continuum or whether it is symptomatic of underlying crucial physical differences between IMTT or Herbig disks and the lower-mass counterparts. The detection of several CO2 isotopologues in XUE 10 highlights the possibility of investigating the isotopic composition of the molecular carriers of the CHNOPS life elements at the protoplanetary disk stage.
@@ Appendix A Local continuum normalization
In order to fit the molecular emission in the observed spectrum, it is first needed to subtract the local dust continuum. We therefore fit and subtracted a spline function in each wavelength window of interest selecting line-free regions, as performed for example by Grant et al. (2023). We evaluated the spectral uncertainty as the standard deviation of the flux in a line-free region close to each fit range. The continuum points used to determine the spline in all windows and the estimated spectral uncertainty are reported in Figure 10. In the 12.9-17.6 m wavelength region with detected emission from the CO2 isotopologues, these species are so abundant to create a pseudo-continuum on top of the dust continuum, corresponding to their strongest emitting features, i.e., fundamental -branch and hot bandheads, particularly those of 12CO2 and 13CO2. This makes it exceptionally challenging to subtract the local dust continuum. We therefore chose the continuum points based on templates of each isotopologue emission for different values of temperature, column density and equivalent emitting radius overlaid on the observed spectrum. We then continuum-subtracted the observed spectrum before fitting each species, selecting line-free regions or regions where the contamination from the other isotopologues was minimal. We note that in Figure 10 it is only shown the interpolated continuum in the original spectrum between 12.9–17.6 m before 12CO2 fit and subtraction.
@@
The only gas phase molecular species capable of creating such a strong feature is diacetylene (C4H2), as shown in Figure 11. In order to reproduce this feature, the gas slab modeling approach requires C4H2 to emit with a very high column density of 3.3 1019 cm-2 at 221 K, and an equivalent emitting radius of 3.3 au ( = 8.1; Figure 13). However, the optically thin line emission expected to arise on top of the optically thick pseudo-continuum is not reproduced in the spectrum, nor the C4H2 line emission which should be visible at 14–14.5 m. The overlap of an optically thick and an optically thin slab model component does not solve the issue.
One may expect to see not only this species, but a ?forest? of hydrocarbonic molecules, starting from the more common acetylene (C2H2) species. This has been observed in the carbon-rich inner regions of some disks around very low-mass stars (e.g., Tabone et al. 2023; Arabhavi et al. 2024; Long et al. 2025; Kanwar et al. 2024). The emission of C4H2 may also be produced by ongoing thermal or UV photo-destruction of carbonaceous grains in the surface layer of XUE 10 (e.g., Visser et al. 2007; Kress et al. 2010; Anderson et al. 2017).
Another possibility may be the emission of PAHs or dust aggregates never seen before in IMTT or Herbig disks at Spitzer resolution (e.g., Meeus et al. 2001; Keller et al. 2008; Valegård et al. 2021). We tested the hypothesis of dust aggregates by fitting the JWST mid-IR dust continuum with the DuCK component of the Dust Continuum Kit with Line emission from Gas (DuCKLinG) modeling tool (Kaeufer et al. 2024). We input the available dust grain opacities for Am Mg-olivine, Am Mg-pyroxene, silica, enstatite and forsterite of 0.1, 1.0, 2.0, 3.0, 4.0, and 5.0 m in size. To model the stellar component, we gave in input the stellar parameters derived for XUE 10 (see Section 3.2). For the disk, we assumed standard parameter priors for the rim, the midplane and the surface, compatible with modeling ranges and observations of Herbig disks. Nonetheless, the ?bump? is not reproduced by any combination of these common solid-state dust species. A more detailed analysis of the dust component and its mineralogy is beyond the scope of this paper, but it will be the subject of a forthcoming work.
@@ Appendix C Spectral classification
The adopted grid spans over 5000 eff 12 000 K, in effective temperature, 2.0 log(g) 6.0 in surface gravity, and -2.0 [Fe/H] +1.0 in metallicity. In order to control for dependencies from the luminosity class over the grid, we first normalized the spectra. To do so, we created a simple Python routine in which a spectrum is split in two parts, over 7730–8400 and 8400–9480 respectively, before and after the Paschen jump. Each part is individually normalized by subtracting a high-order polynomial, and then the two chunks are stitched together into the final normalized spectrum. We note that this routine has been written and optimized for the sole purpose of this analysis.
We proceeded by comparing the normalized (telluric-corrected) FORS2 spectrum with normalized PHOENIX templates, using the reduced -test of the form
@@
In order to evaluate the uncertainty on the fit quantities, we performed an MCMC simulation using the Python emcee package (Foreman-Mackey et al. 2013). We set up uniform priors in the adopted PHOENIX grid for the three fit parameters (eff, log(g), [Fe/H]). We chose the initial positions of the MCMC walkers as the estimated chi-square best-fit values. We imposed the size of the random ?kick? of the walkers from their initial position according to the dynamic range of each parameter space i.e., 500 K in temperature, and 0.5 for surface gravity and metallicity, to ensure a good parameter space sampling.
@@
We run the simulation with 300 walkers and 100 000 steps, adjusted to minimize the autocorrelation times, and increase the number of effective samples in a reasonable amount of computational time. From the posterior distribution of samples, we discarded the first 20 000 samples (20% of the number of steps), and kept a sample from each walker every 20th step, i.e., thinning the posterior distribution by 20. The corner plot of the simulation is shown in Figure 12. The 1 and 2 2D contours are well-defined around the initial positions (blue square). The statistical uncertainties () on the fit quantities are reported on top of the diagonal histograms. The median parameter values of the posterior distribution are 7000 K (eff), 4.0 (log(g)), and +0.5 ([Fe/H]), in agreement with the chi-square fit within error bars.
@@
According to our analysis, XUE 10 can be classified as a F0-F1 type PMS (Gray & Corbally 2009). We combined the spectral information derived with PHOENIX with multi-band photometry to assess visual extinction (V), age, stellar mass (⋆), and bolometric luminosity (bol) of XUE 10. The photometric data include PAN-STARRS, SDSS, Gaia, APASS, and HST in the optical range; DENIS, 2MASS, VISTA, and UKIRT in the mid-IR, and Spitzer and WISE in the far-IR range. Assuming a total-to-selective extinction factor V = 3.3 (Massi et al. 2015; Russeil et al. 2017; Fouesneau et al. 2022), we estimate with VOSA a visual extinction of V 4.8. This result is consistent with the V range of 4.6–4.8 derived using the XUE 10 Gaia color (bp-rp) where a range of spectral types A7–F1 is assumed. We then compared the PHOENIX best-fit template with PMS isochrones of varying age (0.7, 1, 3, 5, 10 Myr) at solar metallicity and at the constrained V, using the PARSEC database (Bressan et al. 2012)666http://stev.oapd.inaf.it/cgi-bin/cmd. We note that PARSEC isochrones are computed by default for an V = 3.1. However, the inferred uncertainties on eff and log(g) are too large to reliably locate XUE 10 on a specific isochrone in this range. XUE 10 is most likely consistent with a 1–3 Myr old PMS in NGC 6357. Starting from the VVV VISTA apparent -band magnitude J = 12.004 0.001, we derive a bolometric luminosity bol = 69 5 ⊙, adopting the bolometric correction in -band for a F1 PMS of Pecaut & Mamajek (2013), the constrained V ( 0.2 from Gaia), and V = 3.3. Considering a PARSEC isochrone of 3 Myr extincted by a factor 4.8 in visual magnitude, the stellar counterpart of XUE 10 is constrained to a mass of 2.5–3.0 ⊙.
@@ Appendix D Fitting tests
Assuming LTE conditions, the fitting procedure consists in comparing a grid of synthetic spectra computed for different values of (gas, tot, em) with the observed continuum-subtracted spectrum through the reduced chi-square test of the form of Eq. 1 (Tabone et al. 2023). Here, obs is the number of MIRI resolution elements covering species features in each fit wavelength window, obs and mod are respectively the observed and the model fluxes to compare, and is the flux uncertainty estimated in a line-free region closed to the spectral region where the species is modeled (see Figure 10). This approach has been extensively used for the spectral fitting of many disk sources so far (see, e.g., Grant et al. 2023; Perotti et al. 2023; Gasman et al. 2023; Ramírez-Tannus et al. 2023; Schwarz et al. 2024; Temmink et al. 2024a), yielding best-fit parameters which are also consistent with the outcome of more sophisticated approaches, e.g., ProDiMo thermo-chemical models (Woitke et al. 2024b) and MCMC simulations (Temmink et al. 2024a). For all molecules, we run a grid of models varying tot (log10-space) in steps of 0.2, gas in steps of 20 K, and em (log10-space) in steps of 0.09. The dynamic ranges explored for each parameter are in Table 4. Given the limited spectral resolution of the MIRI MRS instrument (Jones et al. 2023), we convolved the model spectra with a resolution = 3550 at 4.9–5.2 m, = 3355 at 6.4–7.6 m, = 2600 between 12.9 and 17.6 m, and = 1730 at 18–25 m, where we fit the different species. Then, we resampled the convolved models on the wavelength grid of the observations.
@@ Appendix E Reduced chi-square maps
The maps for all fit species are shown in Figure 13 in order of increasing wavelength in the spectrum. The results for 13CO2, 16O12C18O, and 16O12C17O are reported assuming the best-fit emitting radius found for 12CO2 (em = 1.15 au), and leaving em as a free parameter (see Section 4.2).